A random variable has PDF p_i = 2(1/3)^i and takes on values in the positive integers. What is the probability that it takes on a value between 10 and 20, inclusive? Show your work step by step.


1
Expert's answer
2022-04-06T07:28:08-0400

"P\\left( 10\\leqslant X\\leqslant 20 \\right) =\\sum_{n=10}^{20}{P\\left( X=n \\right)}=\\\\=\\sum_{n=10}^{20}{2\\cdot \\left( \\frac{1}{3} \\right) ^n}=\\left[ \\begin{array}{c}\tgeometric\\,\\,progression\\\\\twith\\,\\,11 terms, b_0=\\left( \\frac{1}{3} \\right) ^{10},q=\\frac{1}{3}\\\\\\end{array} \\right] =\\\\=2\\frac{\\left( \\frac{1}{3} \\right) ^{10}\\left( 1-\\left( \\frac{1}{3} \\right) ^{12} \\right)}{1-\\frac{1}{3}}=\\frac{1}{3^9}\\left( 1-\\frac{1}{3^{12}} \\right) =5.08052\\times 10^{-5}"